a certain force has magnitude of 200N, its x component is -80N. what is its y component and its direction? two possible answers exist ... reply quickly..

Grade:11

1 Answers

vikas askiitian expert
13 years ago

F = 200

Fcos@ = Fx    (x compponent ))        ...............1

Fsin@ = Fy      ( y component )         .............2

from 1

200cos@ = -80

 cos@ = -2/5

here @ can be in 2nd or 3rd quadrant ,

 

if @ is in 2nd quadrant then cos@ is -ve, sin@ is +ve@

cos@ = -2/5 , sin@ = (1-cos2@)1/2 = (21)1/2/5

now , Fy = F sin@ = 200(21)1/2/5 = 184N

 

if @ is in 3rd quadrant then cos@ is -ve , sin@ is -ve

cos@ = -2/5 , sin@ = -(21)1/2/5

Fy = Fsin@ = -184N

 

therefore two possible answers are 184N & -184N

 

aprove if u like my ans